for the given function f(x) 2x+3/1-x find the slope of the tangent line at point x=2


1
Expert's answer
2021-07-16T03:14:25-0400

1.


"f(x)=2x+\dfrac{3}{1-x}"

"f'(x)=(2x+\dfrac{3}{1-x})'=2+\dfrac{3}{(1-x)^2}"

"slope=f'(2)=2+\dfrac{3}{(1-2)^2}=5"

"slope=5"

2.


"f(x)=\dfrac{2x+3}{1-x}"

"f'(x)=(\dfrac{2x+3}{1-x})'=\dfrac{2(1-x)+(2x+3)}{(1-x)^2}"

"=\dfrac{5}{(1-x)^2}"

"slope=f'(2)=\dfrac{5}{(1-2)^2}=5"

"slope=5"